# Implicit In-context Learning

### v1

@@ Abstract
In-context Learning (ICL) empowers large language models (LLMs) to adapt to unseen tasks during inference by prefixing a few demonstration examples prior to test queries. Despite its versatility, ICL incurs substantial computational and memory overheads compared to zero-shot learning and is susceptible to the selection and order of demonstration examples. In this work, we introduce Implicit In-context Learning (I2CL), an innovative paradigm that addresses the challenges associated with traditional ICL by absorbing demonstration examples within the activation space. I2CL first generates a condensed vector representation, namely a context vector, from the demonstration examples. It then integrates the context vector during inference by injecting a linear combination of the context vector and query activations into the model’s residual streams. Empirical evaluation on nine real-world tasks across three model architectures demonstrates that I2CL achieves few-shot performance with zero-shot cost and exhibits robustness against the variation of demonstration examples. Furthermore, I2CL facilitates a novel representation of “task-ids”, enhancing task similarity detection and enabling effective transfer learning. We provide a comprehensive analysis of I2CL, offering deeper insights into its mechanisms and broader implications for ICL. The source code is available at: https://github.com/LzVv123456/I2CL.
@@ 1 Introduction
In-context Learning (ICL) has emerged as a prominent capability of large language models (LLMs) [Brown et al., 2020]. It enables swift adaptation to new tasks during inference by prefixing a few demonstration examples prior to a test query [Wei et al., 2022, Dong et al., 2023]. ICL, characterized by its adaptability and flexibility, has prompted extensive research efforts aimed at optimizing these demonstration examples, or prompts [Rubin et al., 2022, Sorensen et al., 2022, Wu et al., 2023, Min et al., 2022a, inter alia], as well as mitigating their sensitivity to formatting, order, and recency bias [Zhao et al., 2021, Lu et al., 2022, Hao et al., 2022, inter alia].
@@
Despite these advancements, the predominant focus on manipulating demonstration examples within the token space—where tokens are prepended to the query—presents notable limitations. This approach not only escalates computation and memory demands with each additional token but also remains susceptible to the selection and the order of demonstration examples. Such constraints pose significant challenges in scenarios with limited computational and memory resources or variable demonstration profiles, affecting ICL’s scalability and practical utility in real-world applications.
@@
In this study, we explore the usage of demonstration examples within the activation space to surmount these obstacles. Utilizing a decoder-only architecture, we observe that the primary burdens of ICL arise from the computationally intensive multi-head mechanism, which fuses information between demonstration and query tokens, and the memory-intensive key-value caching scheme necessary for retaining demonstration contexts [^1] Without applying key-value cache, one needs to repetitively forward the same demonstration examples.. These observations lead us to explore two pivotal questions: Is there a memory-efficient representation of demonstration examples? And, can we integrate information in a more computationally economical manner?
Figure 1: (Upper-left is better) Comparison of accuracy, inference time and cached memory of different methods on Llama2-7b.
@@
Our findings suggest that both objectives are attainable by first condensing demonstration examples into a compact vector representation and then reintegrating their functionalities within the model’s activation space. Specifically, instead of concatenating demonstration examples into a text sequence before the test query, we independently extract a demonstration vector from each example. These vectors are then aggregated in a permutation-invariant manner to form a unified context vector. During inference, we inject a linear combination of this context vector and the query activations back into the model’s residual streams. We term this scheme Implicit In-context Learning (I2CL), underscoring the absence of an explicit causal relation between the demonstration examples and the test query in the token space.
@@
I2CL offers several merits over standard ICL methods. By condensing demonstration examples into a unified context vector, I2CL requires to cache only a fixed amount of activation vectors, regardless of the length of demonstration tokens. I2CL also maintains a zero-shot inference speed by merging information between demonstration examples and queries through simple linear operations. We evaluate I2CL across three open-source LLMs on nine real-world text classification tasks, where it significantly surpasses the zero-shot and other comparable baselines and achieves results on par with few-shot learning (Figure 1). Importantly, I2CL demonstrates robustness against the variability of demonstration examples. It also facilitates a natural representation of ‘task-ids”, which can effectively indicate task similarity and support transfer learning. The main contributions of this work are summarized as follows:
  1. 1.
    We introduce I2CL, a simple and novel framework that effectively integrates a minimal set of demonstration examples within the activation space. By decomposing conventional ICL into two stages: context vectorization and context-vector injection, I2CL attains few-shot performances at zero-shot costs.
  2. 2.
    We empirically confirm the robustness of I2CL against variations in demonstration examples and its ability to produce a natural representation of task-ids. Leveraging these task-ids, we further propose a transfer learning strategy that can enhance performance on new tasks based on existing ones.
  3. 3.
    We conduct an extensive analysis to thoroughly examine the scheme of I2CL, thereby shedding light on its internal mechanisms and contributing to a deeper understanding of ICL. We also provide a deliberated discussion on the limitations and future directions of I2CL.
2 Related Work
Understanding In-context Learning Besides enhancing ICL, the exploration of ICL’s internal mechanisms has attracted significant research attention. Akyürek et al. [2023] draw parallels between ICL and gradient descent in linear regression tasks, suggesting a fundamental alignment with classical optimization methods. Complementary perspectives from Von Oswald et al. [2023] and Dai et al. [2023] conceptualize ICL as a form of meta-optimization, further enriching our understanding of its operational basis. Concurrently, Xie et al. [2022] interpret ICL through the lens of implicit Bayesian inference, proposing a probabilistic foundation for the learning process. Wei et al. [2023] and Olsson et al. [2022] respectively attribute ICL’s capabilities to the establishment of input-label correspondences and the identification of so-called induction heads, highlighting the intricate interplay between data representation and model interpretability. Despite these advancements, a comprehensive understanding of ICL’s working mechanism remains elusive. Our work, I2CL, contributes additional insights to this domain by substantiating a novel two-stage scheme, which suggests a potential division of the underlying functionality of ICL. Most recently, Label-as-Anchors [Wang et al., 2023] inspects ICL from an information flow perspective and leverages anchor tokens to perform context compression. Similarly, I2CL also enhances the efficiency of ICL, but distinguishes itself by circumventing the need for caching latents of anchor tokens or employing multi-head attention, thereby reducing the inference cost to that of the zero-shot learning.
Activation/representation Engineering An emerging research field, termed activation/representation engineering, also relates to our study. Recent endeavors by Merullo et al. [2023], Turner et al. [2023], Zou et al. [2023] have unveiled the phenomenon of steering vectors, which can be derived from a positive and negative text pair and used to steer the content generation of LLMs. Steering vectors can also be learned via gradient descent [Subramani et al., 2022]. Liu et al. [2024] applies these insights to bolster LLM safety while Li et al. [2023] explores their utility in eliciting more truthful responses from LLMs. To better understand the inner mechanism of activation engineering, the linear representation hypothesis has been studied and discussed in Li et al. [2021], Hernandez et al. [2024], Park et al. [2023]. Our work, I2CL, further corroborates this hypothesis, aligning with and expanding upon the pioneering works by Mikolov et al. [2013], Meng et al. [2023], Hernandez et al. [2023]. Central to this discourse, the idea of task/function vector [Hendel et al., 2023, Todd et al., 2024] resonates with the core premise of I2CL. Both methodologies extract task/function vectors from the demonstration examples and them to improve zero-shot performance. I2CL stands out not only due to its superior performance, but also through its simplicity, namely, its avoidance of the need for task- or architecture-specific hyperparameter tuning, such as selecting attention heads through causal mediation or determining target layers via extra validation sets. Conceptually, we ask an even simpler question: In practice, is there a more economical method to absorb a tiny amount of demonstration examples than by prefixing them in context?
@@ 3.1 Preliminaries
Residual Stream We adopt the mathematical interpretation from Elhage et al. [2021], viewing the hidden states across layers and at each token position as a residual stream. This perspective views each attention head and multi-layer perceptron (MLP) module as engaging with residual streams, facilitating both the addition and deletion of information within them. At any given layer $l$ and token position $t$, the residual stream ${\bm{r}}_{l}^{t}$ is defined recursively:
@@
where ${\bm{a}}_{l}^{t}$ denotes the integrated output of the multi-head attention (MHA) module, and ${\bm{m}}^{t}_{l}$ signifies the MLP’s output, contingent also on ${\bm{a}}_{l}^{t}$. In this framework, MHA promotes the information fusion across residual streams, whereas the MLP functions akin to an associative memory [Geva et al., 2021, Dai et al., 2022] that retrieves information encapsulated within its weights.
In-context Learning Given an unseen task, ICL assumes the existence of a set of $N$ demonstration examples ${\mathbb{D}}=\{d_{1},d_{2},\dots,d_{N}\}$, each comprising an instructional pair $d_{i}=(x_{i},y_{i})$ that includes an input sentence [^2] Input $x_{i}$ also includes formatting texts like: “Question:”, “Answer Type:”. and its corresponding label. ICL operates by first concatenating the demonstration set ${\mathbb{D}}$ with the test query $x_{q}$, forming an input prompt $p=[{\mathbb{D}},x_{q}]$. The objective is then to predict the correct response $y_{q}$ from a finite set of discrete labels ${\mathbb{C}}$ via $y_{q}=\operatorname*{arg\,max}_{y\in{\mathbb{C}}}P(y\mid p)$.
@@ 3.2 Context Vectorization
To overcome the inefficiencies associated with key-value caching system, we isolate the reasoning process of demonstration examples from that of the query, and introduce an independent vectorization of each demonstration pair $d_{i}$. These vectors are subsequently merged in the activation space. Specifically, we define a function $V$, capable of generating a vector representation for each demonstration example: ${\bm{d}}_{i}=V(d_{i})$. Then, a function $F$ is applied to aggregate these demonstration vectors into a unified context vector ${\bm{v}}=F(\{{\bm{d}}_{i}\}_{i=1}^{N})$. Note that $F$ is designed to be permutation invariant, ensuring a unique vector representation for a given set of demonstration examples.
In our implementation, $V$ consists of a pre-trained tokenizer and a frozen LLM parameterized by its pre-trained weights. The demonstration vectors are extracted from the output activations of both MHA and MLP modules at the end residual stream across all layers:
@@
Here, $e$ denotes the end position within a token sequence responsible for next-token prediction, and $L$ represents the total number of layers. For the aggregation function, we compute an element-wise arithmetic mean of each vector component across the demonstration examples:
@@
The rationale of above designs is twofold. First, we argue that the end residual stream encapsulates the essential information from each example. This is supported by the dynamics of next-token prediction and empirical findings from recent studies [Hendel et al., 2023, Zou et al., 2023, Todd et al., 2024]. Second, inspired by the linear representation hypothesis [Park et al., 2023], we premise various demonstration vectors can undergo linear transformations at different levels of abstraction.
Figure 2: A schematic overview of I2CL, including a single layer for the illustrative purpose. I2CL begins by extracting the output activations of the MHA and MLP modules at the end residual stream of each demonstration example. These activation vectors are then fused in a permutation-invariant manner to form layer-wise context vectors. At inference, a linear combination of context vectors and activations from the test query is added to all residual streams.
@@ 3.3 Context Injection
Having established a unique vector representation, ${\bm{v}}$, for the demonstration examples, I2CL seeks to enhance zero-shot performance by integrating this context information with that of the query. While typical implementations leverage the multi-head attention module for this purpose, I2CL utilizes a simpler, yet effective, linear operation to augment the query activations with context vectors.
Consider ${\bm{r}}_{l}^{t}$, the residual stream at the $t$-th token and $l$-th layer of a test query. Instead of merely adding the output activations from the MHA and MLP to their corresponding residual streams, we inject a linear combination of these activations with context vectors:
@@
Here, $\lambda^{a},\beta^{a},\lambda^{m},\beta^{m}$ are four layer-wise scalars used to adjust the proportion to which the context vectors and query activations are blended. They also affect the amplitude of vectors. By default, we apply this information fusion process to all residual streams of a given query, implemented using the forward hook function. Letting $\lambda=0$ and $\beta=1$ (omitting subscripts) will replicate the original zero-shot inference.
Above context injection method incurs minimal computational overhead, involving only two scalar multiplications and an element-wise addition, a process more efficient than the attention mechanism. As a result, the inference speed of I2CL matches that of standard zero-shot learning. Since a constant number of activations is necessary for caching, I2CL also retains approximately the same memory usage as zero-shot. We refer to Figure 2 for a schematic illustration of the context vectorization and injection procedures.
@@ 3.4 Noisy Self-calibration
To achieve generic and nuanced control over the information fusion process, we propose estimating the linear coefficients ${\bm{c}}=\{\lambda_{l}^{a},\beta_{l}^{a},\lambda_{l}^{m},\beta_{l}^{m}\}_{l=1}^{L}$ using a gradient-based method applied to the demonstration examples. It is worth recapitulating that I2CL does not require any additional data beyond the provided demonstration examples. Concretely, we initialize $\lambda=0.1,\beta=1.0$ to promote a modest initial addition of information, and update these coefficients by minimizing the perplexity of label tokens, represented by the following loss function:
@@
where $P(\cdot)$ denotes the induced probability distribution over the entire vocabulary at the end token position from the last layer. To bolster the robustness and adaptability of the linear coefficients to potential downstream variations, we introduce random Gaussian noises $\bm{\eta}\sim\mathcal{N}(\bm{0},\bm{I})$ into the residual streams during the calibration phase:
@@
Given that only a few linear coefficients (totaling $4L$) are updated, this process is remarkably efficient, consuming 1-2 minutes on a single A100 GPU (40G). Importantly, calibrated linear coefficients are demonstration agnostic—they require calibration only once per task and exhibit excellent generalization ability to unseen demonstration examples (see Section 4.3).
@@ 4 Experiments
In empirical section, we begin by detailing the architectures, tasks, and configurations used in our study, followed by a comparative analysis of I2CL against other relevant techniques. We then delve into the formation of context vectors and the characteristics of the calibrated linear coefficients, demonstrating the robustness of I2CL, as well as identifying the function of calibrated coefficients as task-ids. This section concludes with an extensive ablation study to highlight our rationales.
Models We evaluate I2CL using three open-source decoder-only architectures: GPT2-XL [Radford et al., 2019], GPT-J-6B [Wang and Komatsuzaki, 2021], and Llama2-7b [Touvron et al., 2023]. We selected these models based on their suitability for our computational resources and their range in size from relatively small (1.5B) to large (7B). The primary results, reported in subsequent sections, are based on Llama2-7b, with additional results for other architectures detailed in Appendix C.1. We observed consistent trends across all models.
Tasks Initially, we use the four tasks from Wang et al. [2023], including sentiment analysis: SST-2 [Socher et al., 2013], emotion classification: EmoC [Chatterjee et al., 2019], question classification: TREC [Voorhees and Tice, 2000], and topic classification: AGNews [Zhang et al., 2015]. We further enrich our experiments with five additional datasets, encompassing 5-way sentiment analysis: SST-5 [Socher et al., 2013], movie review classification: MR [Pang and Lee, 2005], 14-way topic classification: DBPedia [Zhang et al., 2015], subjectivity status categorization: Subj [Pang and Lee, 2004], and hate speech detection: hate_speech18 [de Gibert et al., 2018]. We leverage the HuggingFace version of the data [Lhoest et al., 2021] and uniformly sample 500 data points from the development set for evaluation.
Experimental Setup Our experimental setup remains consistent across all tasks unless otherwise specified. For each task, we randomly sample five demonstration examples per class [^4] One exception is DBPedia, where we use only one example per class due to the limitation of GPU memory. following the practice described in Wang et al. [2023] to avoid majority label bias [Zhao et al., 2021] and provide a strong few-shot performance. No instruction is used to describe the task. Input sequences are formed using simple manually designed templates (included in Appendix A). For evaluation, we report the macro-average accuracy across nine tasks, computed under five random seeds. For the calibration process, we optimize linear coefficients for $100$ epochs on the same demonstration set using the AdamW [Loshchilov and Hutter, 2019] optimizer. The learning rate starts at $1\times 10^{-2}$ and anneals to $1\times 10^{-5}$ according to a cosine scheduler. This calibration configuration is applied uniformly across all architectures and tasks without tailoring.

### v2

@@ Abstract
In-context Learning (ICL) empowers large language models (LLMs) to swiftly adapt to unseen tasks at inference-time by prefixing a few demonstration examples before queries. Despite its versatility, ICL incurs substantial computational and memory overheads compared to zero-shot learning and is sensitive to the selection and order of demonstration examples. In this work, we introduce Implicit In-context Learning (I2CL), an innovative paradigm that reduces the inference cost of ICL to that of zero-shot learning with minimal information loss. I2CL operates by first generating a condensed vector representation, namely a context vector, extracted from the demonstration examples. It then conducts an inference-time intervention through injecting a linear combination of the context vector and query activations back into the model’s residual streams. Empirical evaluation on nine real-world tasks across three model architectures demonstrates that I2CL achieves few-shot level performance at zero-shot inference cost, and it exhibits robustness against variations in demonstration examples. Furthermore, I2CL facilitates a novel representation of “task-ids”, enhancing task similarity detection and fostering effective transfer learning. We also perform a comprehensive analysis and ablation study on I2CL, offering deeper insights into its internal mechanisms. Code is available at https://github.com/LzVv123456/I2CL.
@@ 1 Introduction
In-context Learning (ICL) has emerged as a prominent capability of large language models (LLMs) (Brown et al., 2020). It enables a swift inference-time adaptation towards new tasks by prefixing a few demonstration examples prior to the query (Wei et al., 2022; Dong et al., 2023). ICL, characterized by its adaptability, has prompted extensive research efforts aimed at optimizing the selection of demonstration examples, or prompts (Rubin et al., 2022; Sorensen et al., 2022; Wu et al., 2023; Min et al., 2022a, inter alia), as well as mitigating their sensitivity to formatting, order, and recency bias (Zhao et al., 2021; Lu et al., 2022; Hao et al., 2022, inter alia).
@@
Despite existing advances, current approaches predominantly focus on manipulating demonstration examples within the token space, where context tokens are prepended to the query. This practice quadratically escalates the computation and memory demands with each additional token and is known to be sensitive to the selection and order of demonstration examples Zhao et al. (2021); Lu et al. (2022); Dong et al. (2023). Consequently, it is challenging to apply ICL under constrained scenarios where computational and memory resources are scarce, and demonstration profiles are uncontrollable, hindering ICL’s scalability and practical utility.
@@
In this study, we explore to harness the demonstration examples within the activation space, seeking for an efficient and robust alternative for constrained environments. Given a decoder-only architecture, we observe that the primary burdens of ICL arise from the computationally intensive multi-head attention mechanism, which fuses information between demonstration and query tokens, and the memory-intensive key-value caching scheme necessary for retaining contextual information [^1] Without applying key-value cache, one needs to repetitively forward the same demonstration examples.. These observations motivate us to investigate the following two questions: Is there a more abstract representation of demonstration examples? And, can we integrate such information into models without resorting to attention mechanism?
@@
Our findings suggest that both objectives are attainable by condensing demonstration examples into a compact vector representation and reintegrating their functionalities within the model’s activation space. Instead of concatenating demonstration tokens before the query tokens, we independently extract a demonstration vector from each example. These demonstration vectors are aggregated in a permutation-invariant manner to form a unified context vector. During inference, a linear combination of the context vector and the query activations is injected into the model’s residual streams as the substitution of the original output activations. We term above scheme Implicit In-context Learning (I2CL), alluding to the absence of explicit demonstration tokens at querying stage.
Figure 1: Upper-left is better. Comparison of accuracy, inference speed and cached memory of different methods on Llama2-7b.
@@
I2CL offers unprecedented merits. By condensing demonstration examples into a unified vector representation, I2CL needs to cache only a fixed amount of activation vectors, independent to the number of demonstration tokens. I2CL maintains a zero-shot inference speed through merging information between demonstration examples and queries using only linear operators. We evaluate I2CL across three open-source LLMs on nine real-world text classification tasks, where it significantly outperforms zero-shot counterpart and other comparable methods. I2CL achieves results on par with few-shot learning with zero-shot inference cost (see Fig. 1). Importantly, I2CL demonstrates robustness against the variability of demonstration examples, and facilitates a natural representation of task-ids that can effectively signify task similarities and foster transfer learning.
Our main contributions can be summarized in three-fold: (1) We introduce I2CL, a novel framework that effectively integrates a minimal set of demonstration examples within the activation space. By decomposing conventional ICL into two stages: context vectorization and context-vector injection, I2CL attains few-shot level performances at zero-shot inference cost. (2) We empirically validate the robustness of I2CL against the variations (i.e., choices and order) of demonstration examples and uncover a natural generation of task-ids, upon which we further propose a transfer learning strategy that can enhance performance on new tasks based on existing anchors. (3) We conduct a comprehensive analysis and ablation to thoroughly examine each component and design choice of I2CL, thereby shedding light on I2CL’s internal working mechanisms.
@@ 2.1 Preliminaries
Figure 2: A schematic overview of I2CL, including a single layer for illustrative purpose.
Residual Stream We adopt the mathematical interpretation from Elhage et al. (2021), viewing the hidden states across layers and at each token position as a residual stream. This perspective treats each attention head and multi-layer perceptron (MLP) module as read-out and write-in operators that engage with residual streams, facilitating the addition and deletion of information within residual streams. At a given layer $l$ and token position $t$, the residual stream ${\bm{r}}_{l}^{t}$ is defined recursively as:
@@
where ${\bm{a}}_{l}^{t}$ denotes the integrated output of the multi-head attention (MHA) module, and ${\bm{m}}^{t}_{l}$ signifies the MLP’s output, contingent on ${\bm{a}}_{l}^{t}$. Within this framework, MHA promotes the information fusion across residual streams, whereas the MLP functions akin to an associative memory (Geva et al., 2021; Dai et al., 2022) that retrieves information encapsulated within its weights.
In-context Learning Given an unseen task, ICL assumes the existence of a set of $N$ demonstration examples ${\mathbb{D}}=\{d_{1},d_{2},\dots,d_{N}\}$, each comprising an instructional pair $d_{i}=(x_{i},y_{i})$ that includes an input sentence [^2] Input $x_{i}$ also includes formatting texts like: “Question:”, “Answer Type:”. and its corresponding label. ICL operates by first concatenating the demonstration set ${\mathbb{D}}$ with the test query $x_{q}$, forming an input prompt $p=[{\mathbb{D}},x_{q}]$. The objective is then to predict the correct response $y_{q}$ from a finite set of discrete labels ${\mathbb{C}}$ via $y_{q}=\operatorname*{arg\,max}_{y\in{\mathbb{C}}}P(y\mid p)$. We acknowledge the versatility and broader implication of ICL, especially when connecting to the generic definition of prompt engineering. In this work, we consider the standard few-shot classification task as our testbed and focus on this scenario in the following sections.
@@ 2.2 Context Vectorization
To overcome the inefficiencies associated with key-value caching system, we isolate the reasoning process of demonstration examples from that of the query, and introduce an independent vectorization process for each demonstration pair $d_{i}$. Extracted vectors are merged in the activation space. Specifically, we define a function $V$, of generating a vector representation for each demonstration example: ${\bm{d}}_{i}=V(d_{i})$, followed by a function $F$ applied to aggregate extracted demonstration vectors into a unified context vector ${\bm{v}}=F({{\bm{d}}_{i}}_{i=1}^{N})$. Note that $F$ is designed to be permutation invariant, ensuring a unique vector representation for a given set of demonstration examples.
In our implementation, function $V$ is parameterized by a pre-trained LLM and its corresponding tokenizer, and operates by collecting output activations of both MHA and MLP modules across layers at the position of end residual stream:
@@
Here, $e$ denotes the last token position that is responsible for next-token prediction, and $L$ represents the total number of layers. For the aggregation function, we compute an element-wise arithmetic mean of each vector component across the demonstration examples:
@@
The rationale of these designs is twofold. First, we argue that the end residual stream encapsulates the essential information from each example. This is supported by the dynamics of next-token prediction and empirical findings from recent studies (Hendel et al., 2023; Zou et al., 2023; Todd et al., 2024). Second, inspired by the linear representation hypothesis (Park et al., 2023), we premise various demonstration vectors can undergo linear transformations at different levels of abstraction.
@@ 2.3 Context Injection
Having established a unique vector representation, ${\bm{v}}$, for the demonstration set, I2CL seeks to enhance zero-shot performance by integrating the contextual information. While conventional methods leverage the multi-head attention module for this purpose, I2CL utilizes a simpler, yet effective, linear operation to augment the query activations with context vectors.
Taking residual stream ${\bm{r}}_{l}^{t}$ as an instance, instead of directly adding the output activations from the MHA and MLP to ${\bm{r}}_{l}^{t}$, we inject a linear combination of these activations with context vectors:
@@
Here, $\lambda^{a},\beta^{a},\lambda^{m},\beta^{m}$ are four layer-wise scalars used to adjust the proportion to which the context vectors and query activations are blended. By default, we apply this information fusion process to all residual streams of a given query. Letting $\lambda=0$ and $\beta=1$ (omitting subscripts) replicates the original zero-shot inference process. Note that it is critical to extract and aggregate information at both MHA and MLP modules (see Section 3.4), suggesting a distinct yet complementary functionality of MHA and MLP.
The proposed context injection method incurs minimal computational overhead, involving only two scalar multiplications and an element-wise addition, a process more efficient than the attention mechanism. As a result, the inference speed of I2CL matches that of standard zero-shot learning. We refer to Figure 2 for a schematic illustration of the context vectorization and injection procedures.
@@ 2.4 Noisy Self-calibration
Thus far, we have presented how to vectorize and reintegrate the contextual information in an efficient and attention-free manner. Herein, we elaborate on how to configure the scale of linear coefficients. It is not uncommon to set those weight scalars as hyper-parameters (Turner et al., 2023; Liu et al., 2024), and manually adjust them for each task in a trial and error fashion. In order to achieve an adaptive and nuanced control over the information fusion process without human-in-the-loop, we propose estimating the linear coefficients ${\bm{c}}=\{\lambda_{l}^{a},\beta_{l}^{a},\lambda_{l}^{m},\beta_{l}^{m}\}_{l=1}^{L}$ using a gradient-based optimization applied to the same set of demonstration examples. Our estimation of linear coefficients does not require any additional data beyond the given few-shot examples.
Concretely, we initialize $\lambda=0.1,\beta=1.0$ to promote a modest initial addition of information, and update these coefficients by minimizing the perplexity of label tokens:
@@
where $P(\cdot)$ denotes the induced probability distribution over the entire vocabulary at the end token position from the last layer. To bolster the robustness and adaptability of estimated linear coefficients to potential downstream variations, we perturb the residual streams with Gaussian noises $\bm{\eta}\sim\mathcal{N}(\bm{0},\bm{I})$ during the calibration phase:
@@
Given above formulations, only a few linear coefficients (totaling $4L$) are updated during the calibration phase, rendering this process remarkably efficient (consuming 1-2 minutes on a single A100 40G). Critically, the calibrated linear coefficients are demonstration agnostic—they require calibration only once per task and exhibit excellent generalization ability to unseen demonstration examples (see Section 3.3). Moreover, these linear coefficients, though are lightweight, can function as effective task-ids that fostering task similarity detection (Fig. 5) and transfer learning (Table 3).
@@ 3 Experiments
In this empirical section, we begin by detailing the architectures, tasks, and configurations used in our study, followed by a comparative analysis of I2CL against other relevant techniques. We then delve into the formation of context vectors and the characteristics of the calibrated linear coefficients, demonstrating the robustness of I2CL, as well as identifying the function of calibrated coefficients as task-ids. This section concludes with an extensive ablation study to underscore the inner working mechanism of I2CL. We refer readers to Appendix C for additional experiments and analysis.
Models We evaluate I2CL on three open-source architectures: GPT2-XL (Radford et al., 2019), GPT-J-6B (Wang & Komatsuzaki, 2021), and Llama2-7b (Touvron et al., 2023). We selected these models based on their suitability for our computational resources and their range in size from relatively small (1.5B) to large (7B). We report results under Llama2-7b in the main content and defer results of other architectures to Appendix C.1. Consistent trends are observed across all architectures.
Tasks We first take the four tasks used in Wang et al. (2023), including sentiment analysis: SST-2 (Socher et al., 2013), emotion classification: EmoC (Chatterjee et al., 2019), question classification: TREC (Voorhees & Tice, 2000), and topic classification: AGNews (Zhang et al., 2015). We then enrich our experiments with five additional datasets, encompassing 5-way sentiment analysis: SST-5 (Socher et al., 2013), movie review classification: MR (Pang & Lee, 2005), 14-way topic classification: DBPedia (Zhang et al., 2015), subjectivity status categorization: Subj (Pang & Lee, 2004), and hate speech detection: hate_speech18 (de Gibert et al., 2018). We employ the HuggingFace version of the data (Lhoest et al., 2021) and sample 500 data points from the validation/test set for evaluation.
Experimental Setup The following configuration is applied to all experiments unless otherwise specified. For each task, we randomly sample five demonstration examples per class [^4] One exception is DBPedia, where we use only one example per class due to the limitation of GPU memory. following the practice described in Wang et al. (2023) to avoid majority label bias (Zhao et al., 2021), and provide a fairly strong few-shot performance. No instruction is further applied to describe the task. Input sequences are formed using simple manually designed templates (included in Appendix A). For evaluation, we report the macro-average accuracy across nine tasks, computed under five random seeds. For the calibration process, we optimize linear coefficients for $100$ epochs on the same demonstration set using the AdamW (Loshchilov & Hutter, 2019) optimizer. The learning rate starts at $1\times 10^{-2}$ and anneals to $1\times 10^{-5}$ according to a cosine scheduler. This calibration profile is applied uniformly across all architectures and tasks without tailoring.